# VINP: Variational Bayesian Inference with Neural Speech Prior for Joint ASR-Effective Speech Dereverberation and Blind RIR Identification

### v2

@@ IV-D1 Speech Dereverberation
Speech dereverberation performance is evaluated in terms of both perception quality and ASR accuracy. We use the commonly used speech quality metrics including perceptual evaluation of speech quality (PESQ) [62], extended short-time objective intelligibility (ESTOI) [63], and deep noise suppression mean opinion score (DNSMOS) [64, 65]. Speech utterances are normalized by their maximum absolute value before evaluation. Higher scores indicate better speech quality and intelligibility.
To demonstrate the effectiveness of our method on ASR systems, we utilize the pre-trained Whisper [66] ‘tiny’ model (with 39 M parameters), ’small’ model (with 244 M parameters), and ‘medium’ model (with 769 M parameters) for ASR evaluation. No additional dataset-specific finetuning or retraining is applied before ASR inference. The word error rate (WER) is used as the evaluation metric. Lower WER indicates better ASR performance.
@@ IV-D2 Blind RIR Identification
We use the accuracy of RT60 and DRR estimation to evaluate the performance of blind RIR identification. We present the mean absolute error (MAE) and the root mean square error (RMSE) of RT60 and DRR between their estimates and ground-truth values over the SimACE test set. Lower MAE and RMSE indicate better results.
@@ IV-E1 Speech Dereverberation
Setting the number of VEM iterations to 100, the dereverberation results are presented in Table II, where the bold font denotes the best results. The results with negative dereverberation effects are marked with a darkened background.
@@
As the metrics of unprocessed speech and oracle speech indicate, a larger ASR system implies a stronger recognition ability for oracle speech and stronger robustness to reverberation. Compared with using a large ASR system alone, introducing an ASR-effective front-end speech dereverberation system can significantly improve the performance of ASR with fewer parameters and lower computational costs. Among the comparison methods, only a few approaches are proven to be effective for ASR, such as GWPE, CMGAN, oSpatialNet*, and the proposed VINP-TCN+SA+S and VINP-oSpatialNet. The classical GWPE yields a marginal improvement in ASR performance. Among the DL-based methods, SkipConvNet, StoRM and TCN+SA+S encounter failures in ASR. TCN+SA+S and StoRM demonstrate marginal effectiveness solely when the ’tiny’ ASR model is employed. The reason for such ASR performance lies in the artifact induced by DNN, which is a common phenomenon in similar DL-based single-channel systems [26, 27]. Different from the comparison methods, VINP employs a linear CTF filtering process to model the reverberation effect and leverages the DNN output as a prior distribution of anechoic source speech. During the VBI procedure, VINP takes into account both this prior information and the observed data. By doing this, VINP circumvents the direct utilization of DNN output and thereby reduces the artifacts. Meanwhile, VINP still utilizes the powerful non-linear modeling ability of DNN. Consequently, VINP exhibits a remarkable superiority over the other approaches in ASR performance. When comparing VINP-TCN+SA+S with its backbone network TCN+SA+S, a gap in WER can be discerned, indicating that VINP can make an ASR-ineffective DNN effective. Both oSpatialNet* and VINP-oSpatialNet achieve ASR performance close to that of oracle speech on REVERB SimData when using the small and medium ASR systems. In other situations, a gap still exists, indicating that VINP can make an ASR-effective DNN even more effective. Because VINP-oSpatialNet provides an anechoic speech prior with higher quality, its ASR performance is superior to VINP-TCN+SA+S. Notably, VINP-oSpatialNet achieves SOTA performance in ASR.
Regarding speech quality, the ESTOI of VINP is roughly the same as that of the backbone network, while the PESQ shows a slight decrease compared to the backbone networks. This indicates that VINP is restricted by the capabilities of the backbone DNN. It is also worth noting that the PESQ increases with the growth of the number of VEM iterations, which we will demonstrate in detail in the experiments presented later. In subjective evaluation, the DNSMOS metrics of VINP-TCN+SA+S and VINP-oSpatialNet reach an advanced level. However, it does not rank the highest among all methods, especially when considering DNSMOS P.835. We hold the view that DNSMOS exhibits a certain preference for enhanced speech. As evidence, it can be seen that the DNSMOS of CMGAN and StoRM is even higher than that of oracle speech, indicating a preference for generative DNNs. In addition, VINP-TCN+SA+S and VINP-oSpatialNet exhibit worse and comparable performance to TCN+SA+S in DNSMOS P.835. However, in our auditory perception, the output speech of VINP exhibits a higher degree of naturalness. To illustrate the characteristics of different methods, examples of the reverberant speech, the output of TCN+SA+S, the output of VINP-TCN+SA+S, and the oracle speech are shown in Fig. 5. Due to the insufficient modeling capabilities of DNN, the output of TCN+SA+S has some overly smoothed artifacts. On the contrary, VINP-TCN+SA+S not only refers to the priors provided by DNNs, but also directly utilizes the observation during the VBI procedure, resulting in a better estimation of the anechoic speech spectrum. Moreover, VINP-TCN+SA+S demonstrates a better performance of background noise control. A series of enhanced audio examples are available online999https://github.com/Audio-WestlakeU/VINP.
@@
To illustrate the influence of VBI iterations on the performance of VINP, we show the average logarithmic likelihood of the complete data (where the constant is omitted) and some metrics on REVERB SimData as the iteration count rises in Fig. 6. As the iteration count increases, the logarithmic likelihood and all metrics shift in a positive direction and tend to converge. Specifically, ESTOI, DNSMOS, and WER exhibit rapid convergence. In contrast, PESQ shows a continuous improvement that closely resembles the behavior of the logarithmic likelihood. This phenomenon indicates that better dereverberation performance can be achieved at the cost of computational complexity. Additionally, the capability of the backbone network determines the upper bound of performance of PESQ, ESTOI and DNSMOS. The logarithmic likelihood of VINP-TCN+SA+S and VINP-oSpatialNet converge to nearly the same values, yet their metrics are different, indicating that when the prior quality varies, the logarithmic likelihood cannot be directly used as an indicator to measure the enhancement performance.
@@ IV-E2 Blind RIR Identification
Also setting the number of VEM iterations to 100, the RT60 and DRR estimation results are illustrated in Table III, where the bold font represents the best results.
VINP surpasses all comparison methods, reaching the SOTA level for all metrics. Theoretically, one limitation of our method is that a CTF filter with a finite length can only model a RIR with a finite length. Nevertheless, it remains sufficient for RT60 and DRR estimation even under large-reverberation conditions, because the signal energy at the tail of the RIR is relatively negligible. From the perspective of spatial acoustic parameter estimation, VINP can estimate RIR well.
@@
BUDDy [35] | 0.089 | 0.132 | 3.93 | 4.57 |
VINP-TCN+SA+S (prop.) | 0.079 | 0.094 | 3.83 | 4.27 |
VINP-oSpatialNet (prop.) | 0.079 | 0.098 | 3.87 | 4.32 |
We also show the MAE of RT60 and DRR estimation as the iteration count rises in Fig. 7. As can be observed, as the iteration count increases, the errors of RT60 and DRR estimation decrease, which indicates an improved quality of RIR estimation. After 30 iterations, VINP-TCN+SA+S and VINP-oSpatialNet start to outperform BUDDy. When the iteration count is less than 70, VINP-oSpatialNet outperforms VINP-TCN+SA+S. Conversely, when the iteration count is larger than 70, VINP-TCN+SA+S is slightly superior to VINP-oSpatialNet. This phenomenon may be related to the DNN structure. Owing to the complexity of the entire system, we are currently unable to draw a conclusion about the preference for a specific DNN architecture for RIR identification. Unlike the experimental results of dereverberation, the identification of RIR does not seem to be sensitive to the capability of DNN when there are enough iterations.
@@ V Conclusion
In this paper, we propose a variational Bayesian inference framework with neural speech prior for joint ASR-effective speech dereverberation and blind RIR identification. By combining the prior distribution of anechoic speech predicted by an arbitrary discriminative dereverberation DNN with the reverberant recording, VINP employs VBI to solve the proposed CTF-based probabilistic graphical model and further estimate the anechoic speech and RIR. The usage of VBI avoids the direct utilization of DNN output but still utilizes its powerful nonlinear modeling capability, and proves to be effective for ASR without any joint training with the ASR system. Experimental results demonstrate that the proposed method achieves superior or competitive performance against the SOTA approaches in both tasks.

### v3

@@ IV-D1 Speech Dereverberation
Speech dereverberation is evaluated in terms of both perception quality and ASR accuracy.
We use the commonly used speech quality metrics including perceptual evaluation of speech quality (PESQ) [69], extended short-time objective intelligibility (ESTOI) [70], and the overall score of deep noise suppression mean opinion score (DNSMOS) P.808 [71] and P.835 [53]. Because DNSMOS P.835 is scale-variant, speech utterances are normalized by their maximum absolute value before evaluation. Higher scores indicate better speech quality and intelligibility. Moreover, log-spectral distortion (LSD) [72] in dB is employed to quantify the difference between the enhanced spectrogram and the clean one. Before calculating LSD, the clean speech is normalized by its maximum absolute value, while the enhanced speech is scaled to match the power of the clean speech. Lower LSD indicates a better estimation of the spectrogram. However, while these metrics are widely adopted to evaluate speech quality, none of them serves as the gold standard. Therefore, we design a subjective absolute category rating (ACR) listening test on REVERB SimData according to ITU-T Rec. P.808 [73]. In the ACR test, a total of 20 subjects (7 females and 13 males) aged from 20 to 55 years were asked to give overall speech quality scores using five-point category-judgement scales, in which a score of 5 indicates excellent speech quality and 1 indicates poor speech quality. In the listening test, all methods were applied to process the same 10 inputs, which were randomly selected from the noisy reverberant speech in REVERB SimData. The unprocessed and clean utterances were also included to screen the subjects. During the test, the listeners were blind to the methods, and the speech samples were arranged in a random order. MOS is used as the final subjective metric. Two subjects (1 female and 1 male) demonstrating a consistent preference for unprocessed over clean speech were excluded from further analysis.
We utilize the popular pre-trained Whisper [74] ‘tiny’ model (with 39 M parameters), ’small’ model (with 244 M parameters), and ‘medium’ model (with 769 M parameters) for ASR evaluation. No additional dataset-specific finetuning or retraining is applied before ASR inference. WER is used as the evaluation metric. Lower WER indicates better ASR performance. Since Whisper is scale-variant, all utterances are normalized by their maximum absolute value before being processed.
@@ IV-D2 Blind RIR Identification
We use the accuracy of RT60 and DRR estimation to evaluate blind RIR identification. We present MAE and the root mean square error (RMSE) of RT60 and DRR between their estimates and ground-truth values over the SimACE test set. Lower MAE and RMSE indicate better results. Notice that the reference RT60s and DRRs are calculated using Eq. (38) and Eq. (39) to ensure fair comparison.
@@ IV-E1 Speech Dereverberation
Empirically setting the number of VEM iterations to 100, the dereverberation results are presented in Table II, where the bold font denotes the best results. The results with negative effects are marked with a darkened background.
@@
As the metrics of unprocessed speech and clean speech indicate, a larger ASR system implies stronger recognition ability and greater robustness to reverberation. Compared with using a large ASR system alone, introducing an ASR-effective front-end speech dereverberation system can significantly improve the performance of ASR with fewer parameters and lower computational costs. Among the comparison methods, only a few approaches are proven to be effective for ASR, such as GWPE, CMGAN, oSpatialNet*, and the proposed VINP-TCN+SA+S and VINP-oSpatialNet. Among the DL-based methods, SkipConvNet, StoRM, and TCN+SA+S encounter failures in ASR, although some of them are effective on the ’tiny’ model. The reason for such ASR performance lies in the unpredictable artificial errors induced by DNN [29, 30, 75]. While these errors do not significantly affect speech perceptual quality, their impact on back-end speech recognition applications remains uncertain. In contrast, larger ASR models are inherently more robust to noise and reverberation and place greater emphasis on extracting the intrinsic features of speech. For these advanced models, however, the artificial distortions introduced by DNN-based dereverberation approaches become more pronounced, often resulting in a decline in recognition accuracy. Different from the comparison methods, VINP employs a linear CTF filtering process to model the reverberation effect and leverages the DNN output as a prior distribution of anechoic source speech. During the VBI procedure, VINP takes into account both this prior information and the observation. By doing this, VINP bypasses the direct utilization of DNN output and ensures the outputs satisfy the signal model, consequently reducing the artifacts. Meanwhile, VINP still utilizes the powerful non-linear modeling ability of DNN. Consequently, VINP exhibits remarkable superiority over the other approaches in ASR performance. When comparing VINP-TCN+SA+S with its backbone network TCN+SA+S, a gap in WER can be discerned, indicating that VINP can make an ASR-ineffective DNN effective. The WERs of oSpatialNet* and VINP-oSpatialNet on SimData with the ’medium’ Whisper model are the same. However, a gap still exists in other situations, indicating that VINP can make an ASR-effective DNN even more effective. A Z-test is conducted between VINP and the backbones at a significance level of 0.05. Results indicate that the differences in WER are significant, except for VINP-oSpatialNet and oSpatialNet* on the Whisper ’medium’ model. When comparing VINP w/o VEM with the original backbones, it becomes evident that the DNNs adapted to speech prior estimation in the proposed VINP yields a certain degree of performance improvement on ASR. However, such an improvement is still insufficient to make TCN+SA+S effective for ASR tasks. Comparing VINP with VINP w/o VEM, it is clear that VEM further improves the ASR performance. Notably, VINP-oSpatialNet achieves SOTA in ASR.
[…]
@@
To illustrate the influence of VEM iterations on the performance of VINP, we show the average expected logarithmic likelihood of the complete data (where the constant is omitted) and some metrics on REVERB SimData in Fig. 5. As the iteration count increases, the logarithmic likelihood and all metrics shift in a positive direction. ESTOI, DNSMOS and WER tend to converge after 40 iterations. In contrast, the logarithmic likelihood and PESQ show a continuous improvement. This phenomenon indicates that better dereverberation performance can be achieved at the cost of computational complexity. With 100 iterations, the logarithmic likelihood of VINP-TCN+SA+S is higher than that of VINP-oSpatialNet, yet the metrics of VINP-oSpatialNet are better. This indicates that when the DNN architecture varies, logarithmic likelihood cannot serve as an indicator to evaluate the enhancement performance of VINP.
@@ IV-E2 Blind RIR Identification
Empirically setting the number of VEM iterations to 300, the RT60 and DRR estimation results are illustrated in Table III, where the bold font represents the best results.
Results show that VINP reaches the SOTA and advanced levels in RT60 and DRR estimation. Theoretically, one limitation of our method is that a CTF filter with a finite length can only model a RIR with a finite length. Nevertheless, it remains sufficient for RT60 and DRR estimation even under large-reverberation conditions, because the signal energy at the tail of the RIR is relatively negligible. From the perspective of spatial acoustic parameter estimation, VINP can estimate RIRs well.
@@
FiNS [36] | 0.113 | 0.167 | 2.15 | 2.64 |
BUDDy [38] | 0.124 | 0.173 | 3.93 | 4.57 |
VINP-TCN+SA+S (prop.) | 0.089 | 0.124 | 3.26 | 3.76 |
VINP-oSpatialNet (prop.) | 0.103 | 0.154 | 2.40 | 2.88 |
We also show the MAE of RT60 and DRR estimation as the iteration count rises in Fig. 8. As the iteration count increases, the errors of RT60 and DRR estimation decrease, indicating improved RIR identification. It appears that the DRR tends to converge after 100 iterations, while the estimation accuracy of RT60 seems to still improve after 300 steps. Therefore, the maximum number of iterations should be selected according to the objective. VINP-oSpatialNet performs better in DRR estimation, while VINP-TCN+SA+S excels in RT60 estimation, which may be related to the DNN structure. Owing to the complexity of the entire system, we are currently unable to draw a conclusion regarding the preference for a specific DNN architecture in RIR identification.
IV-E3 Noise Robustness
Method | REVERB Noise | CHiME4 Noise | ||
5 | ||||
5 dB | 20 dB | 5 dB | 20 dB | |
Unprocessed | 30.0 | 13.3 | 26.8 | 12.9 |
VINP-TCN+SA+S | 21.7 | 9.0 | - | - |
VINP-oSpatialNet | 14.6 | 8.2 | - | - |
Oracle prior + VEM | 8.3 | 8.0 | 8.6 | 8.1 |
Method | REVERB Noise | CHiME4 Noise | ||
5 | ||||
5 dB | 20 dB | 5 dB | 20 dB | |
Unprocessed | 1.15 | 1.48 | 1.11 | 1.43 |
VINP-TCN+SA+S | 1.79 | 2.52 | - | - |
VINP-oSpatialNet | 1.98 | 2.82 | - | - |
Oracle prior + VEM | 2.47 | 3.11 | 2.28 | 3.05 |
Method | REVERB Noise | CHiME4 Noise | ||
5 | ||||
5 dB | 20 dB | 5 dB | 20 dB | |
VINP-TCN+SA+S | 0.406 | 0.089 | - | - |
VINP-oSpatialNet | 0.438 | 0.103 | - | - |
Oracle prior + VEM | 0.415 | 0.071 | 0.427 | 0.098 |
VINP is designed for applications under relatively simple noise conditions. In this part, we adjust both the SNR and noise type in REVERB SimData and SimACE to demonstrate how the noise level and mismatched non-stationary noise affect the performance of VINP. Specifically, we replace the stationary REVERB noise with that from CHiME4 [76], which was collected from four distinct environments: buses, cafes, pedestrian zones, and street intersections. Most noise recordings in CHiME4 are non-stationary, and the transient noise is included. Since there are already numerous DNN-based speech denoising methods, we do not discuss the influence of non-stationarity noise on the capability of DNNs, but use the oracle prior and focus on the effect of VEM. The result using the oracle prior is the upper bound of the proposed framework. WER using Whisper ’tiny’ model and PESQ are shown in Table IV and Table V, respectively. The MAE of RT60 estimation is shown in Table VI. The results demonstrate that neither VINP-TCN+SA+S nor VINP-oSpatialNet achieves the performance ceiling due to errors in the DNN prediction. As the SNR increases, the performance of both speech dereverberation and RIR identification tasks improves. This is consistent with our expectations, as a higher SNR indicates a more reliable observation in the VEM procedure. Compared to the results with matched stationary noise, the performance of VINP degrades under mismatched non-stationary noise conditions, particularly for the dereverberation task.
@@ V Conclusion
In this paper, we propose a variational Bayesian inference framework with neural speech prior for joint ASR-effective speech dereverberation and blind RIR identification. Combining the neural prior distribution of anechoic speech with the reverberant recording, VINP employs VBI to solve the CTF-based probabilistic signal model and further estimate the anechoic speech and RIR. The usage of VBI avoids the direct utilization of DNN output but still utilizes its powerful nonlinear modeling capability, and proves to be effective for ASR systems without any joint training. Experimental results demonstrate that the proposed method achieves superior or competitive performance against the SOTA approaches in both tasks.